2013 Cruise Updates
 

October 25th 21:00 PDT Update

25-Oct-13

On the last day of the Vector expedition, we have successuly deployed the repaired hydrohone array at 170 m in the Strait of Georgia, recovered autonomous mooring from the Saanich Inlet sill location, and recovered the remaining pig bones from Saanich Inlet.

Read more in Daily Log - October 25, 2013


October 24th 21:00 PDT Update

24-Oct-13

Today we successfully deployed the new forensic experiment, recovered both Benthic Boundary Layer (BBL) platform and the IOS hydrophone array. All operations were in the Strait of Georgia. Also, today it was decided not to redeploy the new hydrophone systems (ONC Innovations) that were recovered two days earlier. The hydrophones will require more time to fine-tune and will be re-deployed in 2014.

Read more in Daily Log - October 24, 2013


October 23rd 21:00 PDT Update

23-Oct-13

Day 3 of the expedition featured servicing and re-deployment of the Delta Dynamics Laboratory (DDL) platform, recovery of the Seismic Liquefaction In-situ Penetrometer (SLIP), and an attempt deploy the next forensic experiment in the Strait of Georgia, which was then postponed to the following day.

Read more in Daily Log - October 23, 2013


October 22nd 21:00 PDT Update

22-Oct-13

Another busy day aboard CCGS Vector has come and gone. The ONC engineering crew succesfully serviced and redeployed the core instruments platform in Saanich Inlet, retrieved forensic experiment platform and brought to shore two ONC Inovations' hydrophone arrays for the on-shore servicing.

Read more in Daily Log - October 22, 2013


October 21st 21:00 PDT Update

21-Oct-13

The first day of the VENUS observatory maintenance cruise aboard the Canadian Coast Guard Ship Vector is now complete. A week-long expedition began with servicing instruments at the Fraser River delta location of the the Strait of Georgia array of VENUS. 

Read more in Daily Log - October 21, 2013


CCGS Vector Expedition Begins

20-Oct-13

Between October 20-27, 2013 Ocean Networks Canada is conducting the last scheduled maintenance expedition of the 2013 calendar year. This cruise's focus is maintenance of the VENUS observatory installations from the Canadian Coast Guard Ship Vector using the CanPac Divers' ROV Oceanic Explorer. The expedition's objectives include:

  • Maintain and clean the core sensors on all VENUS instrument platforms (VIP)
  • Recover, service, and redeploy the Delta Dynamics Laboratory near Sand Heads
  • Recover the Seismic Liquefaction In situ  Penetrometer (SLIP) at the Delta site
  • Recover, service, and redeploy three hydrophone arrays in the Strait of Georgia
  • Recover (from Saanich Inlet, with bones), service, and redeploy (Strait of Georgia) the Forensics Platform, with two new pig carcasses.
  • Recover the Bottom Boundary Layer (BBL) platform from the Central Strait of Georgia site

Check the daily logs and cruise status sections for updates.


September 19th 1630 PDT Update

19-Sep-13

A very successful cruise aboard the R/V Falkor has come to an end. The ship docked in Victoria at Ogden Point around 0830 yesterday morning, and we spent the day off-loading equipment and cleaning the labs. The final day of operations was spent at Tully Canyon where we did two extensive survey dives to examine benthic life in the low-oxygen environment. In the evening, we had a barbecue all together on the upper deck of the ship to celebrate the end of the cruise. Many thanks to all those who contributed to make this expedition possible: Schmidt Ocean Institute, the crew of the R/V Falkor, the ROPOS team, our visiting scientists, and the staff supporting us on shore.


September 16th 1900 PDT Update

16-Sep-13

Operations in Barkley Canyon finished at noon today. The last dive, the 600 m transect, completed successfully after a couple of minor delays overnight. We found quite a bit of kelp which was sampled, in addition to sponges and anemones. Now we are doing a multibeam survey and looking forward to two dives starting at midnight at Tully Canyon, a short steam away across the U.S. border.

Read more at Daily Log - September 16, 2013


September 15th 1000 PDT Update

15-Sep-13

Dive R1655, the 300 and 400 metre deep transects of Barkley Canyon, took place overnight. It was very interesting - especially so for those of us who have seen the Ocean Networks Canada sites in Barkley Canyon many times - because of the rugged topography and prevalence of rocks and pebbles that we don't see further down in the canyon. Kelp found, samples taken, and we are now underway to the 1500 m transect site.

Read more in Daily Log - September 15, 2013


September 14th 1520 Update

14-Sep-13

The tether retermination went very quickly, however before we could get back in the water, a power issue with the telemetry can also needed to be resolved. ROPOS worked through the night, allowing us to start Dive R1564 early this morning. In that dive we completed a 200 m deep transect of the head of Barkley Canyon. We encountered a region where fragile pink sea urchins were incredibly abundant. We did suction samples of the sediment surface, push cores and gathered sea urchins and sea cucumbers for analysis. Dive R1655: 300 m and 400 m transects will be starting in a few hours.

Read more in Daily Log - September 14, 2013


September 13th 0945 Update

13-Sep-13

The first transect at 2000 m resulted in some interesting samples including sea pens, sea cucumbers and an anenome for which we currently have no identification. Unfortunately, last night on recovery there was a kink in the ROV tether which has forced us to pause survery operations to reterminate the umbilical. We hope to be diving again early this afternoon.

Read more in Daily Log - September 13, 2013


September 12th 1345 Update

12-Sep-13

This morning we finished the last instrument/experiment deployment dive of the cruise. Next on the programme are visual and sampling transects at staggered depths across Barkley Canyon to study the relationship between oxygen levels and observable fauna. We are at bottom, just about to start the first transect at 2000 m, below the oxygen minimum zone (Dive 1653).

Read more in Daily Log - September 12, 2013


September 11th 0830 PDT Update

11-Sep-13

Dive 1651 at Barkley Canyon POD 3 has just started. In this dive, we will be deploying the first algae treatment of a sediment enrichment experiment. We will be distributing phytodetritus within the frames that we placed in view of the camera yesterday. While waiting for the treatment to settle, we will survey at Wallyland and return to the Coral Cliff to recover the second ADCP.

Read more in Daily Log - September 11, 2013


September 10th 0530 PDT Update

09-Sep-13

The operations at Clayoquot Slope (formerly OPD 889) went well: we swapped the CTD for a CTD-O2, we exchaned SCIMPI's data logger and recovered the broadband seismometer auxiliary platform. Now, the Falkor has arrived at Barkley Canyon, and we are getting ready for a morning dive to install a sonar at Wallyland, clean the crawler and deploy the first stage of a sediment enrichment experiment at POD 3.  The dive will be starting around 0800.

Read more in Daily Log - September 9, 2013 and Daily Log - September 10, 2013


September 8th 1230 PDT Update

08-Sep-13

We have completed all our work in Saanich Inlet and the Strait of Georgia: push cores have been taken, 2 cliff climbs and a transect were completed, the Buoy Profiling System is connected and organisms have been sampled. We are now heading offshore to Clayoquot Slope where we will first perform a navigation system calibration and followed by instrument operations starting early Monday morning.

Read more in Daily Log - September 8, 2013


September 7th 0720 PDT Update

07-Sep-13

The first dive of Leg 2 of the Falkor cruise has started! We are in Saanich Inlet, starting a visual transect of Patricia Bay. The purpose of this dive is to discover the distribution of flatfish in relation to depth and oxygen levels. We will be collecting live samples of these fish for further analysis in the lab.

Read more in Daily Log - September 7, 2013


R/V Falkor Expedition, Leg 2 Begins

06-Sep-13

Today the R/V Falkor set sail from Ogden Point, kicking-off Leg 2 of the "Open Ocean to Inner Sea expedition". Preliminary tasks include several benthic surveys in Saanich Inlet, along with the physical connection of the new Buoy Profiler System recently installed in Saanich Inlet to the VENUS node in Patricia Bay. 

Read more in Daily Log - September 6, 2013

Read more about the expedition in "ONC and Schmidt Team up for Leading Edge Science Research".


R/V Falkor Expediton, Leg 1 is Complete

31-Aug-13

August 31, 15:00PDT Update

Today the R/V Falkor pulled into the Esquimalt Harbour and the first leg of the ONC-SOI 2013 joint expedition come to a conclusion. We packed up our personal possessions, cleaned the laboratories, and dismantled the scientific gear. By mid-afternoon the aft deck was clear and a truck-load of equipment was return to UVic’s Marine Technology Centre. Tomorrow, the loading of ROPOS begins.


August 30th 2300 Update

30-Aug-13

Prior to our most recent exposure to this special oceanographic system, MVP is most often a reference to Most Valuable Player. It is an award, often handed out annually in a sports league, to a player who has contributed the most in a team’s effort to achieve their goals. The MVP on this cruise was, in both senses, the Moving Vessel Profiler.

Read more in Daily Log - Aug 30, 2013


August 29th 2300PDT Update

29-Aug-13

As the penultimate science day, we are really ticking off the boxes of what needs to be accomplished. Yesterday we did a little run along the coastal front next to Vancouver Island. We then zig-zagged yet again along the shelf break, and finally we did a nice tidal time series over the Zeppelin Bank and across the Tully canyon over a flood tide. The final MVP survey tomorrow will be relatively short (12 hours), before we bang off our last CTD and head for home.

Read more in Daily Log - Aug 29, 2013


August 28th 2300PDT Update

28-Aug-13

Today the edits to the plan were minor, and we have started our final push. The weather is grey, with low clouds, irregular seas, and cool temperatures. The forecast calls for an increase in the SE wind. The MVP remains in the water nearly all day, with periodic checks on the fish to inspect for kelp and wear. Our adaptive sampling has a few days left to map some of the key features one last time.

Read more in Daily Log - Aug 28, 2013


August 27th 2300PDT Update

27-Aug-13

The bumpy sea abated through the night, and by morning, the winds were light and the sea a confused wash of old and dying waves. We steam along at our preferred MVP speed of eight knots, and sample with particular purpose. The MVP is used to find exact locations across the continental shelf front for our final set of CTD casts. We are very pleased with these samples. By the end of the day, we have started a marathon of an MVP survey, very likely our last hurrah. If we never give up, luck will find us.

Read more in Daily Log - Aug 27, 2013


August 26th 2300PDT Update

26-Aug-13

There has been a standing forecast the last couple of days for gale force winds. They started to build last night, and outside of Barkley Sound, we are seeing a steady 30 knots. Both the MVP and CTD seem game at first, so the sampling program should not be affected, but we will take a little more care with deck operations.

Read more in Daily Log - Aug 26, 2013


August 25th 2300PDT Update

25-Aug-13

Today a plan for the remainder of the cruise was prepared. We'll see how long it lasts, events change regularly at sea. But it’s a workable plan and includes the remaining cruise objectives. The adaptive science is relentless and gives me no quarter.

Read more in Daily Log - Aug 25, 2013


August 24th 2300PDT Update

24-Aug-13

The day has been a gentle, long sequence of repeated MVP profiles. Counting up into the hundreds, the ship steams along at eight knots, while the profiler surfaces and sinks, again and again.

Read more in Daily Log - Aug 24, 2013


August 23rd 2300PDT Update

22-Aug-13

We complete the second MVP survey along the centre of Tully Canyon, ending near the NEPTUNE Folger site and Bamfield. We let Stephane Gauthier off at Bamfield, so he can get ready for another cruise. We conduct a multi-beam survey of Folger pinnacle, and start the third MVP survey. The oxygen story is both coming into focus and confirming some recent findings by Bill Crawford about the presence of very low oxygen water in the central shelf west of Vancouver Island. At the end of the day, rather dazed and less confused, we see bubbles.

Read more in Daily Log - Aug 23, 2013


August 22nd 0900 PDT Update

22-Aug-13

Today we completed the second MVP survey along the Tully Canyon thalweg, ending near the NEPTUNE Folger site and Bamfield. We let Stephane Gauthier off at Bamfield.  We conducted a multi-beam survey of Folger pinnacle, and the third MVP survey.

Read more in Daily Log - Aug 22, 2013


August 21st 0900 PDT Update

21-Aug-13

Late in the evening Tuesday we started the first moving vessel profiler (MVP) survey. It was completed this morning, and has given us some food for thought. The middle of the day was devoted to calibrating the ship’s new ultra-short baseline (USBL) transponder, before we started an eagerly anticipated MVP transect up the Tully Canyon.

Read more in Daily Log - Aug 21, 2013


August 20th 0900 PDT Update

20-Aug-13

The day has been a very typical oceanographic sampling day, dominated by “classic hydrography”, which is the measurement of water properties as a function of depth at fixed stations across the ocean. Continuing from yesterday, we completed the initial CTD and Rosette survey, including portions of the Barkley Canyon and La Perouse C lines.

Read more in Daily Log - Aug 20, 2013


August 19th 0800 PDT Update

19-Aug-13

This morning at 0800 we arrived at the entrance to Juan de Fuca Strait. A light fog, but relatively calm seas. The morning briefing on the bridge confirms today’s simple schedule: a final configuration of the MVP and then we start an intense CTD/Rosette survey.

Read more in Daily Log - Aug 19 2013


August 18th 09:00 PDT Update

18-Aug-13

At 09:00 PDT R/V Falkor slipped the dock and headed north out of Nanaimo into the central Strait of Georgia. First step is to work out sampling procedures and calibrate ecosounder.

Read more in Daily Log - Aug 18 2013


August 17th 16:00 PDT Update

17-Aug-13

Still at dock in Nanaimo. It has been a busy day aboard the R/V Falkor. We have been both physically and virtually preparing for our departure at 0900 August 18.

Read more in Daily Log - Aug 17 2013


R/V Falkor Cruise Begins

16-Aug-13

Our fourth major expedition of Summer 2013 will be underway on August 18, embarking from Nanaimo, BC on board the Schmidt Ocean Institute's R/V Falkor. We invite everyone to share in daily expedition operations and adventures, via Wiring the Abyss 2013 and Twitter - #Abyss13.

Read more in Daily Log - Aug 16 2013




R/V Thompson Expedition Complete

27-Jun-13

We have now completed our expedition aboard the R/V Thompson and demobilized from ship. We're looking forward to our next expedition about the R/V Falkor, beginning August 17.


June 25th 00:25 PDT Update

24-Jun-13

We are homeward bound! The R/V Thompson is presently in the Strait of Juan de Fuca, on its way to port in Seattle. Our scheduled ETA is 1100, June 25th, following which trucks will arrive to load our gear, and the ONC staff will return to Victoria via the Clipper. The last stop at Clayoquot Slope (formerly ODP 889) was short, but we nevertheless did 2 dives. We recovered the CSEM instrument, investigated a vent hole and visited the SHRIMPI and SCIMPI instruments.


June 23rd 23:45 PDT Update

23-Jun-13

Dive M0017 has come and gone. We deployed four instruments at Cascadia Basin (formerly ODP 1027), our deepest site at 2660 m. We are now steaming to Clayoquot Slope (formerly ODP 889) to recover the CSEM instrument, investigate a vent hole, and collect data from SCIMPI and Shrimpi instruments.  This will be our last stop of the cruise before returning to port in Seattle.


June 23rd 09:00 PDT Update

23-Jun-13

Endeavour is behind us and we have moved on to Cascadia Basin (formerly ODP 1027). At this site, we will deploy a low-frequency (LF) hydrophone, a bottom pressure recorder (BPR) experiment, an acoustic Doppler current profiler (ADCP) and a conductivity-temperature-depth (CTD) instrument with an oxygen sensor. Dive M0017 has just started. The ROV and the instrument package on the ship's wire are descending and will be on bottom shortly.


June 22nd 17:30 Update

22-Jun-13

The cable spool was recovered this morning and is now on deck. Dive M0016 at Main Endeavour Field is underway. The deployment of the Remote Access water Sampler (RAS) just came to a successful end. Next, we are taking some tubeworm samples, and then recovering the ROV. Next stop, Cascadia Basin (formerly, ODP 1027).


June 22nd 01:00 PDT Update

21-Jun-13

We have moved back to the Endeavour Mothra site to recover the Mothra cable and LOCO cable laying system. Dive M0015 has just begun. This afternoon, we dove at the Endeavour North location (RCM-North) to continue the diagnosis of the communications failure. We determined that plugging the existing cable into a different node port did not bring the site online. After performing several other tests, it is suspected that the cable might be at fault for the problem.


June 21st 07:20 PDT Update

21-Jun-13

At the end of Dive M0011, the ROV tether caught on some rocks around the base of Grotto which necessitated maintenance after recovery to the spooling system. Dives M0012 and M0013 also had to end early due to related problems. The deployment of the Remote Access water Sampler (RAS) was interrupted.  We are taking a break from Main Endeavour Field to continue the diagnosis of the communications issues at the RCM-North site, but will return later to finish MEF operations.


June 20th 07:50 PDT Update

20-Jun-13

Dive M0011 is coming to an end. We need to recover the ROV in order to reset the gas tights for RAS (remote access water sampler) deployment. In the last day, we completed Tempo-Mini's deployment, a visual survey of Grotto for photogrammetry, deployment of temperature loggers near Tempo-Mini and we designed and lowered new bio-buckets for tubeworm collection, which will occur on the next dive. RAS will be lowered on the ship's wire this morning.


June 19th 16:00 PDT Update

19-Jun-13

Dive M0011 has now been underwater for nearly 40 hours. The crew, ROV pilots, science and engineering teams are working hard to keep up with 24 hour operations. Tempo-Mini deployment is ongoing; the ROV is getting ready to position the instrument on Grotto. Today is a beautiful sunny day and we are still fortunate with calm seas.


June 19th 01:10 PDT Update

18-Jun-13

Dive M0011 ongoing and proceeding well. Following the last update, BARS has been successfully deployed and water samples and temperature profiles for COVIS were completed. We also took tubeworm samples from North Tower and gas tights for BARS calibration. At the time of this writing, Tempo-Mini is waiting for deployment on the ship's wire and will be placed next. Other upcoming tasks: temperature logger replacement, photogrammetry survey, short-period seismometer and auxilliary platform deployment, and further sampling.


June 18th 09:20 PDT Update

18-Jun-13

Dive M0011 at Main Endeavour Field in progress. Instruments are being lowered to the seafloor on a specially built platform on the ship's wire and guided to the bottom by the ROV. The first load included the 600 kHz ADCP (Acoustic Doppler Current Profiler) and BARS (Benthic and Resistivisty Sensors) vent probe. Early this morning, the ADCP was deployed. We are currently in the process of placing BARS in a vent on the Grotto crown.


June 17th 19:00 PDT Update

17-Jun-13

We are moving on station to begin dive operations at Main Endeavour Field. In the next two days, we will deploy RAS, BARS, Tempo-Mini, a 600kHz ADCP, temperature loggers, and do biological and visual sampling. We anticipate being on bottom around midnight tonight, PDT, but it will still take us some time to begin operations, so the first deployments will start in the early morning.  We left RCM-North without successfully diagnosing the failure. The spare junction box did not resolve the issue; we will return later in the cruise.

 


June 17th 00:00 PDT Update

16-Jun-13

We have moved on to the RCM-North site to diagnose the communications failure at that location. Dive M0010 is in progress. The first test is to evaluate the functioning of the junction box. We have lowered a spare instrument platform on the ship's wire and we will be attempting to connect it shortly. The attempts to mate the connector to LOCO on previous dives were not successful and therefore we were unable to lay the Mothra-Midpoint cable up to now.


June 16th 14:00 PDT Update

16-Jun-13

Dive M0009 in progress. We attempted to bring LOCO, the cable laying system, up to 250 m in order to try to mate the connector to the cage. We were able to plug it in, but the connector would not stay mated. The ROV is presently on the way back to the bottom with the spool. We will not be able to lay the cable at this time, however, we are in contact with Oceaneering and they are working on a solution. In the meantime, if we can't lay the cable, we intend to go to RC-North to diagnose the failure at that site.


June 16th 00:30 PDT Update

15-Jun-13

Dive M0008 is underway. The goal of this dive is to bring the cable layer (LOCO) attached to the Mothra cable spool up to a shallower depth at which we hope it will be possible to mate the connector between the cage and LOCO. Our tests on Dive M0007 revealed that the dummy connectors could be successfully mated between 250 and 500 m depth. When connected, we will return the spool to the seafloor and begin the Mothra-Midpoint cable lay. Once started (at the earliest, Sunday morning PDT), the lay will take approximately 12 hours to complete.


June 15th 09:40 PDT Update

15-Jun-13

More dive woes overnight. We had to do an emergency recovery because the pan & tilt was damaged on the main HD camera during dive M0006 before the connection to LOCO could be established. Weather is still co-operating. All hoping for good news later today.


June 14th 20:30 PDT Update

14-Jun-13

The ROV is back on deck again. We were unable to start the cable lay because of problems mating the connector to power the cable laying system. The ROV team is working on the issue, and we should be diving again soon.


June 14th 11:30 PDT Update

14-Jun-13

We are beginning the Mothra-Midpoint cable lay. The heading required for the cable lay might interfere with satellite communications from the ship. (There is a small "block zone" for visibility of the satellite with the antenna on the ship.)  If there are outages in the live feed, don't panic and please check back later!


June 14th 09:30 PDT Update

14-Jun-13

Dive M0005 is now underway. The end of the cable has been connected to the Mothra Instrument Platform; cable lay to Midpoint upcoming. The last operation of Dive M0004 was to recover four acoustic releases which will sit between the cage and the spool to provide a release mechanism should the cable lay be interrupted. M0004 completed around 04:30 this morning.


June 14th 00:00 PDT Update

13-Jun-13

Dive M0004 is coming to an end. The spool of cable for the Mothra-Midpoint cable lay has been deployed on the ship's wire and readied by the ROV for the cable lay. We will perform the lay on the next dive overnight and into the morning. Weather is perfect for operations: flat seas.


June 13 09:25 PDT Update

13-Jun-13

Dive M0003 has just started; dive objective to lay Mothra to Midpoint cable. M0002 was another short dive, following which several issues were resolved overnight with the ROV and navigation. The weather and seastate are perfect for cable laying so we are optimistic about upcoming operations.


June 13th 00:15 PDT update

12-Jun-13

We are currently at Mothra running a CTD. A first test dive (M0001) with the ROV revealed some minor problems which got sorted out and we are about to conduct another test before we go on the first science dive (Mothra).


June 12th 11:30 PDT update

12-Jun-13

We are approximately 2 hours away from Endeavour, heading to Mothra for our first operations.  We will start with a CTD cast at the site, and then evaluate weather conditions for starting the Mothra to Midpoint cable lay. Winds have been 15-20 kn this morning with a approximately 2 m seas.


June 11th 15:40PDT Update

11-Jun-13

We are currently underway, steaming through Puget Sound en route to the Strait of Juan de Fuca. From there, we will transit to the Endeavour site where we will start our maintenance operations. The ROV has completed two important tests: a successful wet test at the dock in Seattle overnight, and a navigation test this afternoon. We are taking advantage of the 28 hour transit to finish organizing ourselves onboard for the upcoming operations.


June 10th 10:00PDT Update

10-Jun-13
Mobilization is in progress aboard the R/V Thompson for the Ocean Networks Canada June Cruise. We anticipate being ready to sail Tuesday morning. The Oceaneering Millenium ROV is onboard and nearly ready for wet test. All hands are busy with setup of video transmission, network systems, instrument preparation, and securing the deck for transit. The rest of the science party will be arriving later this afternoon on the Clipper from Victoria. Stay tuned!


June 10th 08:00PDT Update

10-Jun-13
Welcome to the R/V Thompson cruise scheduled for June 11 to 26th, 2013. We are currently mobilizing in Seattle, Washington, with crew and visiting researchers arriving from around the world. This expedition will feature the Oceaneering ROV “Millenium,” operated by Canpac Divers. Check out our live video to see the "Millenium," and stay tuned for updates in this space!


May 27th to June 9th

29-May-13

No Expedition Operations at this time. 
Watch for our Thompson expedition - June 9 to June 26.


May 27th 10:00PDT Update

27-May-13

Our expedition aboard the CCGS Tully has concluded and we are now demobilizing from ship at the Institute of Ocean Sciences dock in North Saanich, BC.


May 25th 08:00PDT Update

25-May-13
We are currently transiting from the Strait of Georgia to Saanich Inlet. We will be diving this morning at 1000 to complete a series of cliff surveys in Squally Reach. This afternoon we will shift to our Saanich Node site for one final dive before coming alongside.


May 24th 09:00PDT Update

24-May-13

This morning we are sitting off the mouth of the Fraser River at our Delta Dynamics Laboratory site. We will be diving this morning to work on a hydrophone array. The outflow from the Fraser plus very large tides give us a very narrow time window in which to complete the dive. This afternoon we will shift to our East Node site for another dive.


May 23rd 07:00PDT Update

23-May-13

This morning we are in Saanich Inlet to recover the Micro Squid platform and replace a damaged oxygen sensor. We will work on the package and then redeploy it shortly before lunch. The Tully will then transit to the Strait of Georgia where we will dive to conduct maintenance on our hydrophone arrays.


May 22nd 16:00PDT Update

22-May-13

We have just completed a successful dive in Tully Canyon. We are now conducting a short sonar survey and will then begin the transit back to Saanich Inlet. This marks the end of the NEPTUNE maintenance operations for Tully Leg #2.

We will be diving at 0730 tomorrow morning in Saanich Inlet to conduct maintenance on the VENUS observatory.


May 22nd 07:00PDT Update

22-May-13

Winds have dropped but the seas are still too high in Barkley Canyon to dive.  We are heading to an alternate site in Tully Canyon where conditions look more favourable.

We should be diving at 0830.  Our plans are to spend the day at Tully Canyon and depart for Saanich Inlet in the late afternoon. 

Tomorrow we will  conduct maintenance dives on the VENUS observatory.  

 


May 21st 08:00PDT Update

21-May-13

Winds have been building steadily overnight and are now blowing from the north at 35 kts. Seas have increased and we have been forced to cancel diving operation for today (see Deck Cam feed). We will complete our CDT casts and continue with sonar surveys.

We will remain at Barkley Canyon in the hope that the weather abates tomorrow morning.


May 20th 14:30 Update

20-May-13

With the successful deployment of the Folger Deep platform, our inshore engineering work is now complete. We are heading offshore to our Barkley Canyon sites to prepare for ROV dives tomorrow morning. On our way to Barkley we will conduct CTD casts and then start a new series of sonar surveys in a location suspected of having gas seeps. Our dives should start at 0730 Tuesday morning, weather permitting.

0730 - Dive at Barkley Hydrates site, sonar tripod positioning and visual survey

1330 - Dive to location of gas seeps detected during sonar survey


May 19th 11:50PDT Update

19-May-13

We have just successfully recovered the VPS platform. The operation was completed very smoothly by the Tully deck crew. The sun is shining out here, but the seas have been building, so the captain has made the decision that we should not dive this afternoon.

Instead, we are going to do some more deep water CTD rosettes and continue the echosounder survey for bubble plumes from last night. The current plan is still to deploy Folger Deep Monday morning.


May 19th 0050PDT Update

18-May-13

We're right on schedule at the time of this update:

Overnight: Echosounder survey (from Tully)

Sunday morning: Dive to recover VPS platform

Sunday afternoon:

  • Barkley Hydrates
    • Deployment of second sonar tripod on winch
    • Dive for visual survey of Wallyland; positioning of tripod; coral cliff survey
  • Echosounder survey (from Tully)


May 18th 0800PDT Update

18-May-13

Here is an update of our operations and plans for Saturday and Sunday:

  • CTD rosettes in Effingham Inlet
  • Science dives in Effingham Inlet with Verena Tunnicliffe
  • Redeployment of Folger Deep
  • Echosounder surveys looking for bubble plumes
  • Potential dives at identified locations.

 


May 17th 1630PDT Update

17-May-13

We are in Barkley Sound, waiting for a crew change to be completed.

  • Ian Kulin, Jonathon Miller, Reece Hasanen and Reyna Jenkyns are disembarking.
  • Adrian Round (as expedition leader), Ian Beliveau and Karen Douglas are joining us.
  • Verena Tunnicliffe (UVic) and Bob McDonald (of CBC's Quirks & Quarks) are spending one night with us during our exploration of Effingham Inlet.

Here is an update of our operations and the plans for the next few days:

Currently:

  • Tully in Barkley Sound; crew change in progress from Bamfield
  • Folger Deep platform on board for maintenance
  • PODs 1, 2, 3, 4, and Barkley Hydrates have been redeployed.

Today and tomorrow:

  • CTD rosettes in Effingham Inlet
  • Science dives in Effingham Inlet
  • Redeployment of Folger Deep


May 17th 0800PDT Update

17-May-13

Currently staging Oceanic Explorer dive OE0060 at Folger Passage. Goals for this dive:

  • Recover hydrophone
  • Disconnect bottom pressure recorder
  • Disconnect Folger Deep instrument platform
  • Recover existing cable to node
  • Prepare Folger Deep instrument platform for recovery to ship




May 15th 2000PDT Update

15-May-13

Dive OE0057 was completed with an early end due to ROV issue.

  • POD 4 reconnected (although not powered because Barkley Hydrates is not yet connected).
  • Unfinished at POD 4 due to early end: CTD on tripod not in place.
  • POD 3 recovered; will be maintained overnight
  • No overnight dive.


May 14th 2130PDT Update

14-May-13

Today, we:

  • Deployed Pod 4
  • Dove to Pod 4 (dive OE56), but ROV lost power and dive was aborted 
  • Deployed one tripod for new sonar installation at Barkley Hydrates via winch
  • Completed maintenance on Barkley Hydrates platform
  • Deployed Barkley Hydrates platform via winch

Tomorrow's plans are to:

  • Recover Plankton Pump (Dive OE57 beginning at ~7:30 am)
  • Connect Pod 4 (Junction Box 05)
  • Deploy CTD tripod 
  • Connect Barkley Hydrates instrument platform
  • Tasks for Kongsberg sonars
    • Locate and inspect winch-deployed tripod 
    • Identify landing zone for second tripod 
    • Survey proposed deployment sites
  • Wally Land tasks 
    • Inspect shell and wood experiments 
    • Recover yellow markers 
    • Perform photomosaic survey
  • Place marker beacon at Pod 3
  • Recover two parking positions
  • Recover Pod 3


May 14th 08:00PDT Update

14-May-13

Barkley Canyon Pod 4 is being deployed from ship. Once deployment is finished, the Oceanic Explorer will be launched for dive OE0056 to connect Pod 4 and prepare Pod 3 for recovery.

Barkley Hydrates Instrument Platform remains on deck for now.


May 13th 18:40PDT Update

13-May-13

Dive OE0055 complete. ROV on deck.


May 13th 14:13PDT Update

13-May-13

Dive OE0055 is on. Diving to Barkely coral cliffs.

Location: Barkley Canyon                                                                        

Objectives

·  Deploy autonomous Nortek profilers (DeviceIDs: 23163,23164)

·  Visual survey of coral cliffs

Full Dive Plan OE0055

 


May 13th 12:40PDT Update

13-May-13

Dive OE0054 is complete. ROV on deck; the Barkley hydrates platform is coming up on board first time since 2009. 


May 13th 08:55PDT Update

13-May-13

Dive OE0054 is on.

Location: Barkely Canyon

Objectives:

·  Place marker beacon

·  Disconnect Wally

·  Disconnect Barkley Hydrates Instrument Pod and clear cables safely away

·  Recover Barkley Hydrates Instrument Pod (DeviceID: 10002)

Full Dive Plan OE0054

 


May 12th 20:15PDT Update

12-May-13

Dive OE0053 is complete. ROV and Instrument Pod (POD) 4 on deck.


May 12th 10:15PDT Update

12-May-13

Dive EO0053 in progress.  

Location: Barkley Canyon

Objectives:

· Connect POD 1 (DeviceID: 10011) to network

· Deploy Dragonfish camera system (DeviceID: 23152, 23074)

· Deploy Nortek Profiler (DeviceID: 11302)

· Deploy Kongsberg (DeviceID: 11301)

· Visual transect survey

· Organism sampling

· Recover bait trap

· Recover marker beacon

Full Dive Plan OE0053


May 12th 06:12PDT Update

12-May-13

Dive OE0052 is complete. ROV on deck.


May 11th 18:45PDT Update

11-May-13

Dive OE0052 is on. Weather remains to be a constraint.

Location: Barkley Canyon 

Objectives (updated @ 19:30):

·         Prepare POD 4 Recovery

o   Visual inspection

o   Visual transect survey

o   Prepare platform for recovery

o   Disconnect POD 4 from network and clear cables safely away

·         Prepare POD 3 Recovery

o   Visual inspection at POD 3

o   Prepare Kongsberg (DeviceID: 22789) for recovery

o   Recover Nortek Profiler (DeviceID: 12003) and secure to IP

o   Recover Sediment Trap (DeviceID: 12005) and secure to IP

o   Disconnect POD 3 from network and clear cables safely away

·         Place marker beacon at POD 1

·         Visual survey from POD 1 to Hydrates IP

·         Inspect shell/wood experiment

·         Survey area around Barkley Hydrates for later maintenance work.

·         Recover POD 4

 

Full Dive Plan OE0052


May 11th 16:30PDT Update

11-May-13

Dive OE0051 Complete.


May 11th 11:15PDT Update

11-May-13

Dive OE0051 is on. ROV is in the water. Live video is streaming.


May 11th 08:00PDT Update

11-May-13

Preparing for dive OE0051. Weather is a constraint today.

Location: Barkely Canyon

Objectives:

·   Connect POD 2 JB-03 (DeviceID: 10018) to network

·   Deploy Sediment Trap (DeviceID: 11008)

·   Deploy Dragonfish camera system (DeviceID: 23153, 23076)

·   Deploy Nortek Profiler (DeviceID: 11203)

·   Deploy Kongsberg (DeviceID: 12007)

·   Visual transect survey

·   Organism sampling

·   Recover bait trap

 

 

Full Dive Plan OE0051

 


May 10th 14:00PDT Update

10-May-13

Preparing for the next Dive – OE0050. Scheduled to start early PM.

Location: Barkley Canyon 

Objectives:

 

·  Secure floats to camera tripod

·  Visual inspection of Instrument Pod 

·  Deploy bait trap

·  Visual star-configuration transect

·  Prepare Kongsberg (DeviceID: 11301) for recovery

·  Recover Nortek Profiler (DeviceID: 11302) and secure to IP

·  Recover camera system (DeviceID: 12170, 12126) and secure to IP

·  Disconnect POD 1 (DeviceID: 10011) from network and clear cables safely away.

 

Full Dive Plan OE0050

For details, see Daily Log 10 May


May 10th 07:00PDT Update

10-May-13

Dive OE0049 dive is scheduled for 7:30am. The plan is to complete remaining tasks from OE0048 dive and replace the CTD for Upper Slope. 

Location: Barkley Canyon

Objectives:

• Deploy CTD tripod (DeviceIDs: 23031, 23047) at Upper Slope

• Visual Upper Slope IP (DeviceID: 10003) Inspection • Recover CTD (DeviceID:22607) at Upper Slope

• Recover Dragonfish camera system (DeviceID: 23072, 12127) and secure to Instrument Pod

• Recover Sediment Trap (DeviceID: 11008) and secure to Instrument Pod

• Inspect VPS Platform (DeviceID: 22612)

• Disconnect POD 2 from network and clear cables safely away. 

Full Dive Plan OE0049


May 9th 14:55PDT Update

09-May-13

Planning the first dive of the Leg 2 - Servicing NEPTUNE observatory. Dive OE0048. 

Location: Barkley Canyon

Dive  Objectives: 

• Visual IP Inspections

• Position marker beacon

• Deploy bait trap

• Visual transect

• Prepare Kongsberg (DeviceID: 12007) for recovery

• Recover Nortek Profiler (DeviceID: 11203) and secure to IP

• Recover Dragonfish camera system (DeviceID: 23072, 12127) and secure to IP

• Recover Sediment Trap (DeviceID: 11008) and secure to IP

• Inspect Vertical Profiler System (VPS) Platform (DeviceID:22612)

• Disconnect POD 2 from network and clear cables safely away

• Recover CTD (DeviceID:22607) at Upper Slope.

Full Dive OE0048 Plan

For details, see Daily Log, 9 May  


May 8th 22:30PDT Update

09-May-13

Leg 2 (May 9-28) of the "Wiring the Abyss 2013" on Tully begins.

The plan for the second leg of the “Wiring the Abyss 2013” aboard CCGS Tully is to service instruments at two locations of the NEPTUNE observatory of Ocean Networks Canada. In particular we will focus on Folger and Barkley Canyon locations off the west coast of Vancouver Island.

The planned service includes:

·  Recover, perform maintenance and redeploy Folger Deep platform

·  Recovery, maintenance and redeployment of four Barkley Canyon pods (instrument platforms) and surrounding instruments

·  Recover and perform maintenance and deploy at Barkley Hydrates

·  Install sonar system at Barkley Hydrates

·  Perform sampling at all Folger and Barkley sites and in Effingham Inlet.

For details, see Daily Log, 8 May


May 8th 13:45PDT Update

08-May-13

The Tully has returned to IOS in Patricia Bay for a science crew exchange and re-mobilization for the second leg of this operations and maintenance expedition. Plans for today include the installation of a new satellite system through Oceaneering. Tomorrow, Leg 2 of the cruise should see us heading out into the Northeast Pacific Ocean to service NEPTUNE observatory sites Barkley Canyon and Folger Passage. Taking the reins from Richard Dewey, we will now be hearing from Expedition Leader Ian Kulin, who is Ocean Networks Canada's Associate Director for Marine Operations.


May 6th 23:30PDT Update

07-May-13

Tomorrow (May 7), Verena Tunnicliffe will join us to complete a two-year research effort. Planned are video survey dives up into Patricia Bay, and at least two vertical video profiles along the steep cliff walls deeper in the Inlet. And then Leg #1 will come to an end on May 8. Leg #2 of the Tully cruise will be devoted to servicing several off-shore sites of the NEPTUNE observatory. 

For details, see Daily Log, 6 May


May 5th 23:50PDT Update

05-May-13

The ship is now steaming towards Saanich Inlet. Over the next few days we will re-assemble the Saanich Inlet Instrument Platform and the Oregon State University-led Benthic Oxygen Flux experiments. Benthic surveys are required for two camera projects, and during the last dive we will recover a camera for maintenance. All going well, we should be back at the Institute of Ocean Science dock for the exchange by early afternoon, May 8.

For details, see Daily Log, 5 May 


May 5th 09:45PDT Update

05-May-13

Currently preparing for a dedicated dive to investigate and sort out the navigation system.


May 5th 00:05PDT Update

04-May-13

Yesterday, the Bottom Boundary Layer platform was deployed and connected to the Strait of Georgia Central Node. Today's plans call for deployment of the Seismic Liquefaction In-situ Penetrometer (SLIP) and Delta Hydrophone Array near the Fraser Delta node, followed by possible deployment of the Eastern Hydrophone Array near the Strait of Georgia Eastern Node.

For details, see Daily Log, 4 May


May 4th 00:20PDT Update

03-May-13

Today we will finish the sonar surveys at 6AM, and steam to our SoG Central Node site for one platform deployment (the Bottom Boundary Layer package) and one platform recovery (the autonomous CTD mooring). During these deck operations, we will assemble the complicated Seismic Liquefaction In-situ Penetrometer (SLIP). It will be deployed during a slack tide in the early afternoon. This leaves most of the day for Tom to re-assess his hydrophone arrays, and deploy them on Sunday.

For details, see the Daily Log, 3 May


May 2nd 22:00PDT Update

03-May-13

Tomorrow we pick up two of Tom Dakin's students (Hannan and Kristen) and Gwyn Lintern of NRCan from Steveston, and continue the deployments near the Delta Node. First up will be the Delta Dynamics Laboratory (DDL) at the end of its own kilometre extension cable, up-slope from the Delta Node. This cable is presently being spooled onto the ROV reel. We will now add an extra dive at the end of the day to recover the ONCCEE platform.

For details, see the Daily Log, 2 May


May 1st 23:00PDT Update

02-May-13

The ship slowly streamed into the Strait of Georgia early this morning. After a few days of strong winds, we brought calm to the Strait. But the mighty Fraser is cranking up it's flow, as the Rocky Mountain snow begins to melt. With a 5 fold increase in fresh water volume in May, there is an equal increase in the suspended sediment. The plume had been detected in our Ferry data to have been broadening over the last few weeks, and it clearly now fills a majority of the southern Strait.

The next few days we be at the Delta Laboratory site. After a staff picking up at Steveston early May 2, we will scout out the northern end of the Delta cable, and get the Delta mini-Node re-built, deployed, and connected. The first of two tech-demo hydrophone deployments may finish the day. On May 3, Gwyn Lintern (NRCan) will join us and we will lay a cable and deploy a new DDL platform. May 4 will focus our efforts on the Seismic Liquefaction In-situ Penetrometer (SLIP). This leaves a final hydrophone array and the Bottom Boundary Layer (Dalhousie and MUN) platform for SoG Central.

For details, see the Daily Log, 1 May


May 1st 08:00 Update

01-May-13

Today we'll focus on the recently (April 26) laid Delta extension cable. Our first dive will connect the south end to the SoG East Node. Another dive at the north end of the extension cable will be necessary to secure it, and determine where to deploy the new Delta mini-Node. Time permitting, we'll deploy and connect the mini-Node. The Fraser River freshet is starting, so we might encounter strong surface currents, laden with silt and sediments.

First Dive, diving soon:

Dive OE0028 Objectives:

  • Find Node and East VIP
  • Head east and find cable
  • Follow south and get cable end, take to Node, connect
  • Head SSE and locate autonomous platform, recover

 


April 30th 23:00 Update

30-Apr-13

We finished rigging the navigation beacon assembly and got away from the dock reasonably early. A multitasking effort to test the ROV navigation system while conducting the first CTD and Rosette cast, accommodated a first dive before lunch. This went smoothly. The afternoon saw another dive and platform recovery and the deployment of a vital autonomous mooring at the entrance to Saanich Inlet. We deployed the new forensics platform after dinner and then dove to connect it. A quick return to the IOS dock to exchange a dozen platforms, and we're now heading over to the Strait of Georgia.

For details, see the Daily Log, 30 April.


April 29th 20:00 Update

29-Apr-13

We will either sail at 06:15 or 07:15 tomorrow morning, depending on status of ROV systems. Plans for tomorrow include:

  • CTD and Rosette at SI VIP site/depth to calibrate oxygen measurements.
  • First dive to recover bones from pigs 7 and 8 and HD webcam-platform. Transfer all gear to forensics platform.
  • Second dive to recover SI VIP.
  • Deploy new forensics platform with pigs 9 and 10.
  • Deploy ROV to connect forensics.

On board today for forensics operations: Dr. Gail Anderson and Chantal Turpin

Report by Dr. Richard Dewey, Expedition Leader

For details, see the Daily Log, 29 April


Change to Twitter

29-Apr-13

On 29 April 2013, our Twitter handle for expedition operations changed from @nc_operations to @oceanetworksops.


Cruise Season Begins

29-Apr-13

On April 29, ONC embarks on a series of maintenance cruises to service both the VENUS and NEPTUNE cabled observatories. Over the next two months, we’ll execute three of our eight 2013 expeditions, including two back-to-back cruises on the CCGS John P. Tully (April-May), servicing facilities first in the Salish Sea (VENUS) and then off-shore at Folger Passage and Barkley Canyon (NEPTUNE). 





RV Falkor - Leg 2 Dive Plans (PDF)


Dive R1660 Tully Canyon
September 17 2013 Tully Canyon

Dive R1659 Tully Canyon
September 16 2013 Tully Canyon

Dive R1658 Tully Canyon
September 16 2013 Tully Canyon

Dive R1657 Barkley visual survey
September 15 2013 Barkley Canyon

Dive R1656 Barkley visual survey
September 15 2013 Barkley Canyon

Dive R1655 Barkley visual survey
September 14 2013 Barkley Canyon

Dive R1654 Barkley visual survey
September 13 2013 Barkley Canyon

Dive R1653 Barkley visual survey
September 12 2013 Barkley Canyon

Dive R1652 Barkley Canyon Deployment
September 11 2013 Barkley Canyon

Dive R1651 Barkley Canyon Deployment
September 11 2013 Barkley Canyon

Dive R1650 Barkley Hydrates Deployment
September 9 2013 Barkley Canyon

Dive R1649 Patricia Bay Transect
September 9 2013 Clayoquot Slope

Dive R1649 Strait of Georgia East Push Cores
September 8 2013 Strait of Georgia East

Dive R1648 Patricia Bay Transect
September 7 2013 Saanich Inlet

Dive R1647 BPS Connect
September 7 2013 Saanich Inlet

Dive R1646 Saanich Inlet Cliffs
September 7 2013 Saanich Inlet

Dive R1645 Patricia Bay Transect
September 7 2013 Saanich Inlet